OpenStudy (anonymous):

Use a graphing utility to solve the equation on the interval 0°< x < 360°. Express the solution(s) rounded to one decimal place. cos^2 x + cos x - 1 = 0

OpenStudy (anonymous):

first find the value of cos(x) by solving the given equation.

OpenStudy (anonymous):

Substitute y = cos x, then y^2 + y -1=0. By quadratic formula, \[y=\frac{-1 \pm \sqrt5}{2}=\cos x\]Only the plus (of the plus or minus) will work, since the absolute value of cos x is never more than one. The arccos function will give you angles that correspond to the value of cos x.

OpenStudy (campbell_st):

go and download geogebra http://www.geogebra.org/cms/ this will allow you to graph the curve... as the question asks. The solutions will be where the curve cuts the x-axis

OpenStudy (anonymous):

im still not understanding

OpenStudy (anonymous):

i already download that

OpenStudy (anonymous):

say y = cosx then your equation will become y^2+y-1=0 graph this equation.

OpenStudy (campbell_st):

ok... the only problem is that it will give the decimal values... and you will need to set the x-axis scale to radians...

OpenStudy (campbell_st):

type your function into geogebra as \[f(x) = (\cos(x)^2 + \cos(x) - 1\] then use the roots function \[roots[f, 0, 2\pi]\] it will show the roots as points... but only as a decimal... an alternative would be wolfram alpha which would probably give exact values. http://www.wolframalpha.com/

OpenStudy (campbell_st):

oopps the function should be \[f(x) = (\cos(x))^2 + \cos(x) - 1\]

OpenStudy (anonymous):

i still not have the solution lol

OpenStudy (anonymous):

i graph it but i dont see any number

OpenStudy (campbell_st):

ok... do you see a any points A, B, C, etc..?

OpenStudy (anonymous):

nope there is no abc

OpenStudy (campbell_st):

hold on...

OpenStudy (campbell_st):

the error in your curve is that you need to enter cos^2(x) as (cos(x))^2 this will give the correct curve

OpenStudy (anonymous):

f(x)=(cos(x))^2+cos(x)−1?

OpenStudy (campbell_st):

here is my graph

OpenStudy (campbell_st):

thats correct...

OpenStudy (campbell_st):

so once you have that you will need to use the roots command for the start and end value of the curve just use zero and 7 so the command is roots[f,0,7] and the points A and B should appear...

OpenStudy (campbell_st):

and then in the algebra window you'll have a decimal approximation.